# Emulators for stellar profiles in binary population modeling

### v1

@@ Abstract
Knowledge about the internal physical structure of stars is crucial to understanding their evolution. The novel binary population synthesis code POSYDON includes a module for interpolating the stellar and binary properties of any system at the end of binary MESA evolution based on a pre-computed set of models. In this work, we present a new emulation method for predicting stellar profiles, i.e., the internal stellar structure along the radial axis, using machine learning techniques. We use principal component analysis for dimensionality reduction and fully-connected feed-forward neural networks for making predictions. We find accuracy to be comparable to that of nearest neighbor approximation, with a strong advantage in terms of memory and storage efficiency. By delivering more information about the evolution of stellar internal structure, these emulators will enable faster simulations of higher physical fidelity with large-scale simulations of binary star population synthesis possible with POSYDON and other population synthesis codes.
@@ 1 Introduction
The study of binary evolution stands as an essential and complex area in astrophysics, offering a unique probe into the underlying physical processes governing the lives of most stars and their impact on their environments. Binary star systems serve as invaluable laboratories for testing theoretical models and advancing our understanding of stellar evolution and the production of compact objects and their rejuvenation as electromagnetic and gravitational-wave sources. Furthermore, a large fraction of high-mass stars are born in binaries, making binarity a key factor in massive star evolution (e.g., Duchêne and Kraus, 2013; Chini et al., 2013). A standard way to study the diversity of binary stars and their evolutionary paths is with population synthesis codes, which enable simulation of populations of stellar binaries [^1] Although population synthesis methods have been used for electromagnetic sources for almost 40 years, the recent discoveries of compact binary gravitational-wave sources have led to renewed interest.. This has been especially useful in recent years with LIGO-Virgo-KAGRA opening the door to a significant sample of gravitational-wave discoveries of compact object mergers (Abbott et al., 2019, 2021, 2023). In comparing observed distributions of compact object merger properties like masses, spins, and redshifts, to simulated evolution outcomes of synthetic binary populations, we can better constrain uncertain stages like common envelope evolution and supernovae.
@@
There is a broad slate of existing binary population synthesis codes, most of which rely on analytical fits of single-star evolution tracks based on Hurley et al. (2000), which fit to stellar evolution models computed by Pols et al. (1998). These include SeBa (Portegies Zwart and Verbunt, 1996), binary_c (Izzard et al., 2004, 2006, 2009), StarTrack (Belczynski et al., 2008), COSMIC (Breivik et al., 2020), COMPAS (Riley et al., 2022), and MOBSE (Giacobbo et al., 2023). This approach is very computationally efficient as no actual binary evolution models are used and it is therefore widely used for population-level studies. However, it relies on a very old, outdated set of single-star, non-rotating models and require a large number of additional recipes to account for all the effects of binary evolution. These approximations can lead to conclusions that are inconsistent with population studies based on evolutionary models that self-consistently simulate evolution of the star’s structure and binary interactions (e.g., Gallegos-Garcia et al., 2021; Klencki et al., 2021; Patton and Sukhbold, 2020). Some codes aim to improve model accuracy by interpolating detailed single-star evolutionary tracks, which involves self-consistent treatment of wind mass-loss, from look-up tables and adding basic binary prescriptions, e.g., COMBINE (Kruckow et al., 2018) and SEVN (Spera et al., 2019), or by self-consistently applying effects of mass transfer to single-star evolution but only fully computing the evolution of one star at a time for computational constraints, e.g., BPASS (Eldridge and Stanway, 2009). Most recently Fragos et al. (2023) introduced POSYDON, a binary population synthesis code that represents the current state-of-the-art in fully self-consistently solving stellar structure equations for both stars concurrently with orbital evolution using the 1D stellar structure code MESA (Fragos et al., 2023).
@@
The detailed stellar evolution simulations done with MESA are computationally expensive, with each binary evolutionary track sometimes taking over 100 CPU hours to run (Fragos et al., 2023), so low-cost machine learning models trained on smaller (tens of thousands of binaries) pre-computed MESA grids are needed to generate large populations of binaries. These machine learning models emulate outcomes of MESA evolution (Fragos et al., 2023). The first version of POSYDON includes a newly developed initial-final interpolation module for for mass-transferring binaries at all evolutionary stages, which provides values for scalar stellar and binary properties including mass, effective temperature, radius, and orbital period. At the end of each evolutionary stage, POSYDON also saves information about the full internal stellar structure of the final model generated as part of the MESA simulation. According to convention in stellar astrophysics, we use the term “stellar profiles” to refer to the dependence of any physical quantity that varies along the interior of the star.
@@
Stellar profiles contain valuable information that is not available from integrated or surface quantities —information which can be used to constrain later evolutionary stages and to investigate anomalous behavior. For example, the density profile of a star at the end of a MESA evolutionary stage can be used to inform the following core-collapse stage. 3D simulations have shown that the density profile of a star plays a major role in explodability. Early accretion through a large, sharp drop in the density profile can cause the confining ram pressure to abruptly decrease and the shock to rapidly expand, resulting in an explosion (Vartanyan et al., 2018, 2019, 2021). The outcome of this explosion has proven to be very sensitive to the location of the density jump (Boccioli et al., 2023).
@@
The density and angular rotation profiles of a progenitor can be used to estimate remnant properties. Using the mass and angular momentum distributions of the star, which can be calculated from these profiles, Bavera et al. (2020) add infalling shells of material to the black hole and associated accretion disk in a process described by Batta and Ramirez-Ruiz (2019). The mass and angular momentum added from this material can be used to estimate the mass and spin of the resulting black hole throughout and after collapse.
@@
The chemical composition profiles can be used to calculate the core-envelope boundary, which can then be used to determine whether the system has enough energy to eject a common envelope, a critical phase for the formation of interacting binaries with compact objects (Dewi and Tauris, 2000; Ivanova, 2011; Kruckow et al., 2016). The angular rotation profile can be used to calculate the spins of resulting black holes, whose distributions can be compared to observational black hole spin distributions (Bavera et al., 2020). Having accurate emulated profiles readily available will open the door to many lines of investigation, especially when it comes to custom user-provided prescriptions for uncertain evolutionary stages. This is especially useful for cases where a user wants to test out a prescription on the fly without being limited by the POSYDON default or having to generate an entirely new set of grids.
@@
Due to the correlations between the profile points in the stellar profile dataset, profiles cannot be interpolated in the same way that the global quantities like stellar mass are in the existing interpolation module of POSYDON, which is described in Section 3.1 of this paper. For that reason, we present here a new methodology to emulate stellar profiles, which uses machine learning techniques to both reduce the dimensionality and provide full predictions of stellar profiles for key physical quantities.
@@ 2 Data
The properties of the three POSYDON binary grids are listed in Table 1, and more explanation is provided in Fragos et al. (2023). Grid values for initial primary masses, compact object masses, and orbital periods are spaced logarithmically, whereas initial binary mass ratio values for the HMS-HMS grid are evenly distributed in linear space [^3] This distribution is somewhat modified for the CO-HMS grid during post-processing because the “initial” binary parameters are replaced with the binary parameters at the onset of Roche-lobe overflow, as described in Fragos et al. (2023).. While the v1 grids capture evolution at solar metallicity, POSYDON version 2 will include grids for 8 discrete metallicities: $2Z_{\odot}$, $Z_{\odot}$, $0.45Z_{\odot}$, $0.2Z_{\odot}$, $10^{-1}Z_{\odot}$, $10^{-2}Z_{\odot}$, $10^{-3}Z_{\odot}$, $10^{-4}Z_{\odot}$ (Andrews et al., in prep.).
To evaluate the performance of interpolators trained on each of these MESA grids, we have three corresponding testing grids of binaries that are generated by random sampling of the same parameter spaces as the main binary grids. The outputs of these MESA grids include various properties of the binary at the end of MESA evolution, such as the binary orbital period, structural status of the stars, mass transfer status, and integrated or surface quantities of the stars, such as mass, radius, effective temperature, and luminosity. POSYDON also saves the stellar profiles at the end of MESA evolution, which show the structure of quantities like log density, log temperature, and hydrogen mass fraction at discrete sampling points throughout the star from the center to the surface as a function of either radius or mass. The placement of profile points is dictated by the adaptive mesh in the MESA model; points are more densely sampled in regions where higher resolution is necessary to characterize the complex behavior exhibited in the star (Paxton et al., 2011). The specific quantities saved after a MESA run can be specified by the user generating their own grids; the default settings are explained in detail in Fragos et al. (2023). As an example of the profiles generated by MESA, in Figure 1 we show the final log density profiles of the binaries in the CO-HMS testing grid.
@@ 2.1 Pre-processing
$\varepsilon_{i}=|\frac{\rho_{200}(m_{enc,i})-\rho_{400}(m_{enc,i})}{\rho_{400}(m_{enc,i})}|,$ (1)
@@ 3 Methods
Neural networks are a powerful and readily available tool commonly used to learn complex patterns and relationships in large datasets (e.g., Watt et al., 2020). We use fully-connected neural networks because of their superior ability to approximate functions, and because they evaluate very quickly after training, which is essential for large-scale applications of function approximation problems such as in binary population synthesis. While they are less interpretable than, for example, tree-based methods, the high accuracy, scalability and computational efficiency of evaluation provide a good fit for our profile emulation task. Neural network emulation of astrophysical simulations have previously been employed by Ribas et al. (2020), who use fully-connected neural networks to emulate synthetic protoplanetary spectral energy distributions at any point in their model’s parameter space. In this application, neural networks make it possible to efficiently sample posterior distributions for inference on protoplanetary disk masses, which would not be possible with the full radiative transport modeling. Auld et al. (2008) likewise use fully-connected neural networks to emulate cosmic microwave background power spectra to ultimately conduct parameter estimation on cosmological parameters.
Gaussian process regression is sometimes used in emulation of astrophysical models, for example stellar spectra (Czekala et al., 2015) and cosmological simulations (Habib et al., 2007). However, Gaussian processes are not well suited for high-dimensional problems or large datasets due to the large amount of memory required in the optimization process (Rasmussen and Williams, 2006).
@@
To reduce our profiles’ dimensionality we use principal component analysis (PCA), which is a common technique in astrophysics. PCA provides a representation space in which data points are most widely distributed by choosing axes along which the dataset displays the most variance. The first principal component is along the direction with maximal variance, the second principal component is orthogonal to the first along the direction that maximizes the residual variance in the dataset, and so on (Ivezić et al., 2014).
@@ 3.1 Emulating density profiles
Figure 2: Principal components of CO-HMS grid density profiles plus offsets for visual separation. The first eight components are shown plotted in order of explained variance from top to bottom.
@@ 3.1.1 Dimensionality reduction
We first reduce the dimensionality of our normalized profiles to simplify the task assigned to the neural network. The smoothly-varying nature of stellar quantities like density and temperature make PCA a suitable method. Using PCA, we map our 200-dimensional density profiles (shown in Figure 1) onto a lower-dimensional latent space spanned by a set of orthogonal 200-dimensional vectors, called principal components, along which we can maximally express the variance in the training set. The first eight principal components of density profiles from the CO-HMS grid are shown in Figure 2. Each training profile is projected onto this set of principal components to obtain the projection weights, which are used to create a reconstruction of the original profile as a weighted sum of the principal components.
@@ 3.1.2 Network Architecture
Each model in our emulator is a fully-connected feed-forward neural network, with the input layer consisting of three input nodes (for each of the dimensions of the MESA grids, which are the two stellar masses and orbital period). The networks have 8 output nodes for the profiles’ principal component weights. The depth and width of the network were chosen to balance model complexity and prediction accuracy with computational efficiency. The optimal configuration for our network has 12 hidden layers of 256 nodes each, as shown in Figure 5. We experimented with including normalization layers (such as batch normalization) to stabilize learning and potentially improve convergence, but we found that this led to worse performance, potentially having interfered with the model’s ability to capture sufficient complexity. As a result, the normalization layers were excluded from the final architecture.
@@
We find that separating our data according to mass transfer class improves performance, so for each grid we produce one machine learning model for each mass transfer class. This is consistent with the existing initial-final interpolators from POSYDON v1 (Fragos et al., 2023). For training each of these profile emulation models, we used an 80/20 split to produce a training set and validation set. This step prevents overfitting and ensures the model’s ability to generalize outside the training data. Each network was initialized randomly, and given up to 3000 epochs to train to ensure the model had enough time to converge. However, to further prevent overtraining we used an “early stopping” feature to stop training if the validation loss does not improve for 100 epochs. This number, called the “patience”, was determined through experimentation to give the best results in optimization.
@@
The models were optimized based on a mean squared error (MSE) loss function placing equal emphasis on each component’s weight. It may seem more intuitive to base the loss function on the accuracy of profile predictions rather than the weights, because the accuracy of the weights is incidental to the ultimate goal of profile emulation. However, in practice we found that using a loss function based on the reconstructed profiles rather than the component weights ultimately made a minimal difference to the prediction accuracy. Hence, we chose the MSE loss of the weights for its faster evaluation time. We chose the Adam optimization algorithm, which is a stochastic gradient descent method that is very flexible (Kingma and Ba, 2014). We found Adam to give better results than SGD (Ruder, 2016), RMSprop (Hinton, 2012), AdamW (Loshchilov and Hutter, 2019), and Adagrad (Duchi et al., 2011).
@@
We use the ReLU (Rectified Linear Unit) activation function for all of the hidden layers, which allows the networks to learn non-linear behavior in the training set. We found ReLU to give better results than other activation functions including tanh (hyperbolic tangent) and sigmoid, which we found constrained layer outputs too tightly and reduced the potential complexity afforded by additional layers. The output layer has linear activation to ensure that final predictions for principal component weights are unconstrained. The model was implemented using TensorFlow (Abadi et al., 2015) using the Keras API (Chollet et al., 2015).
@@
Figure 6: Flowchart showing procedure for emulating density profiles. Red boxes are model components from POSYDON’s existing scalar interpolation module, blue boxes are the model components presented in this work, and white boxes are data products. POSYDON’s scalar interpolation uses the initial parameters and associated MESA grids to predict final scalar properties at the end of an evolutionary stage. For density profiles, the final mass and central log density of each star is predicted by this interpolator; the emulator presented here then produces a density profile which is re-scaled by these predicted values.
@@
The overall procedure for obtaining predictions (illustrated in the flowchart in Figure 6) is as follows: we perform PCA on the training profiles to obtain a set of eight principal components and the eight principal component weights corresponding to each training profile. We then train neural networks with the training binaries’ initial conditions (star 1 mass, star 2 mass, orbital period) as input and eight principal component weights as outputs. With this network in place, we predict the weights for the profiles of a star in any new binary, which combined with the principal components construct predicted normalized profiles. We then renormalize these profiles using values for central log density and final star mass predicted with POSYDON’s scalar interpolation module, which predicts these quantities from the initial conditions and the associated MESA grid. With this configuration and the design choices detailed above, we achieve sufficiently good performance for profile emulation. Future work experimenting with different choices or architectures may further improve accuracy.

### v2

@@ Abstract
Knowledge about the internal physical structure of stars is crucial to understanding their evolution. The novel binary population synthesis code POSYDON includes a module for interpolating the stellar and binary properties of any system at the end of binary MESA evolution based on a pre-computed set of models. In this work, we present a new emulation method for predicting stellar profiles, i.e., the internal stellar structure along the radial axis, using machine learning techniques. We use principal component analysis for dimensionality reduction and fully-connected feed-forward neural networks for making predictions. We find accuracy to be comparable to that of nearest neighbor approximation, with a strong advantage in terms of memory and storage efficiency. By providing a versatile framework for modeling stellar internal structure, the emulation method presented here will enable faster simulations of higher physical fidelity, offering a foundation for a wide range of large-scale population studies of stellar and binary evolution.
@@ 1 Introduction
The study of binary evolution stands as an essential and complex area in astrophysics, offering a unique probe into the underlying physical processes governing the lives of most stars and their impact on their environments. Binary star systems serve as invaluable laboratories for testing theoretical models and advancing our understanding of stellar evolution and the production of compact objects and their transformation into electromagnetic and gravitational-wave sources. Furthermore, a large fraction of high-mass stars are born in binaries, making binarity a key factor in massive star evolution (e.g., Duchêne and Kraus, 2013; Chini et al., 2013). A standard way to study the diversity of binary stars and their evolutionary paths is with population synthesis codes, which enable simulation of populations of stellar binaries [^1] Although population synthesis methods have been used for electromagnetic sources for almost 40 years, the recent discoveries of compact binary gravitational-wave sources have led to renewed interest.. This has been especially useful in recent years with LIGO-Virgo-KAGRA opening the door to a significant sample of gravitational-wave discoveries of compact object mergers (Abbott et al., 2019, 2021, 2023). In comparing observed distributions of compact object merger properties like masses, spins, and redshifts, to simulated evolution outcomes of synthetic binary populations, we can better constrain uncertain stages like common envelope evolution and supernovae.
@@
There is a broad slate of existing binary population synthesis codes, most of which rely on analytical fits of single-star evolution tracks based on Hurley et al. (2000), which fit to stellar evolution models computed by Pols et al. (1998). These include SeBa (Portegies Zwart and Verbunt, 1996), binary_c (Izzard et al., 2004, 2006, 2009), StarTrack (Belczynski et al., 2008), COSMIC (Breivik et al., 2020), COMPAS (Riley et al., 2022), and MOBSE (Giacobbo et al., 2023). This approach is very computationally efficient as no actual binary evolution models are used and it is therefore widely used for population-level studies. However, it relies on a very old, outdated set of single-star, non-rotating models and require a large number of additional recipes to account for all the effects of binary evolution. These approximations can lead to conclusions that are inconsistent with population studies based on evolutionary models that self-consistently simulate evolution of the star’s structure and binary interactions (e.g., Gallegos-Garcia et al., 2021; Klencki et al., 2021; Patton and Sukhbold, 2020). More recent codes allow for additional flexibility in the adopted stellar physics by providing different choices of single-star evolutionary tracks as alternatives to those used in Hurley et al. (2000). This has been done by using look-up tables and adding basic binary prescriptions, for example in COMBINE (Kruckow et al., 2018) and SEVN (Spera et al., 2019). Another approach is to self-consistently apply effects of mass transfer to single-star evolution, but only fully compute the evolution of one star at a time for computational constraints, as BPASS does (Eldridge and Stanway, 2009). Most recently Fragos et al. (2023) introduced POSYDON, a binary population synthesis code that represents the current state-of-the-art in fully self-consistently solving stellar structure equations for both stars concurrently with orbital evolution using the 1D stellar structure code MESA (Paxton et al., 2010).
@@
The detailed stellar evolution simulations done with MESA are computationally expensive, with each binary evolutionary track sometimes taking over 100 CPU hours to run (Fragos et al., 2023). Using low-cost machine learning models trained on smaller (tens of thousands of binaries) pre-computed MESA grids is one of the ways to efficiently generate large populations of binaries. These machine learning models emulate outcomes of MESA evolution (Fragos et al., 2023). The first version of POSYDON includes a newly developed interpolation module for mass-transferring binaries at all evolutionary stages, which provides final values for scalar stellar and binary properties including mass, effective temperature, radius, and orbital period as a function of initial binary parameters of stellar mass and orbital period. At the end of each evolutionary stage, POSYDON also saves information about the full internal stellar structure of the final model generated as part of the MESA simulation. According to convention in stellar astrophysics, we use the term “stellar profiles” to refer to the dependence of any physical quantity that varies along the interior of the star.
@@
Stellar profiles contain valuable information that is not available from integrated or surface quantities —information which can be used to constrain later evolutionary stages and to investigate anomalous behavior. For example, the density profile of a star can be used to inform its potential core-collapse stage. 3D simulations have shown that the density profile of a star plays a major role in explodability. Early accretion through a large, sharp drop in the density profile can cause the confining ram pressure to abruptly decrease and the shock to rapidly expand, resulting in an explosion (Vartanyan et al., 2018, 2019, 2021). The outcome of this explosion has proven to be very sensitive to the location of the density jump (Boccioli et al., 2023).
@@
The density and angular rotation profiles of a progenitor can be used to estimate remnant properties. Using the mass and angular momentum distributions of the star, which can be calculated from these profiles, Bavera et al. (2020) add infalling shells of material to the black hole and associated accretion disk in a process described by Batta and Ramirez-Ruiz (2019). The mass and angular momentum added from this material can be used to estimate the mass and spin of the resulting black hole throughout and after collapse. The distributions of predicted black hole spins can then be compared to observational distributions (Bavera et al., 2020).
The chemical composition profiles can be used to calculate the core-envelope boundary, which can then be used to determine whether the system has enough energy to eject a common envelope, a critical phase for the formation of interacting binaries with compact objects (Dewi and Tauris, 2000; Ivanova, 2011; Kruckow et al., 2016).
@@
Having accurate emulated profiles readily available will open the door to many lines of investigation in binary population synthesis studies, especially when it comes to custom user-provided prescriptions for uncertain evolutionary stages. This is especially useful for cases where a user wants to test out a prescription on the fly without being limited by default settings or having to generate an entirely new set of grids.
@@
Due to the correlations between the profile points in the stellar profile dataset, profiles cannot be interpolated in the same way that the global quantities like stellar mass are in the existing interpolation module of POSYDON, which is described in Section 3.1 of this paper. For that reason, we present here a new methodology to emulate stellar profiles, which uses machine learning techniques to both reduce the dimensionality and provide full predictions of stellar profiles for key physical quantities. This is a flexible approach which aims to provide a starting point for efficient emulation of stellar structure. While demonstrated here with density profiles from our POSYDON simulations, our framework can be applied to any physical property of interest.
For the predictions, we leverage neural networks, a powerful and readily available tool commonly employed to learn complex patterns and relationships in large datasets (e.g., Watt et al., 2020). Specifically, we utilize fully-connected neural networks due to their superior ability to approximate functions and because they evaluate very quickly after training, which is essential for large-scale applications of function approximation problems such as in binary population synthesis. Previous work has successfully used neural networks to emulate stellar evolution simulation outputs, including single-star evolutionary tracks (Maltsev et al., 2024), fundamental stellar properties of $\delta$ Scuti stars (Scutt et al., 2023), and observable stellar properties (Lyttle et al., 2021). More broadly, the use of machine learning methods in general to emulate various aspects of stellar evolution has grown in recent years, including the use of normalizing flows to emulate single-star evolutionary grids (Hon et al., 2024), augmenting such grids with Gaussian process regression (Li et al., 2022), and mapping observable quantities of stars to fundamental stellar parameters using random forests (Bellinger et al., 2016).
@@ 2 Data
The properties of the three POSYDON binary grids are listed in Table 1, and more explanation is provided in Fragos et al. (2023). Grid values for initial primary masses, compact object masses, and orbital periods are spaced logarithmically, whereas initial binary mass ratio values for the HMS-HMS grid are evenly distributed in linear space [^3] This distribution is somewhat modified for the CO-HMS grid during post-processing because the “initial” binary parameters are replaced with the binary parameters at the onset of Roche-lobe overflow, as described in Fragos et al. (2023).. While the v1 grids capture evolution at solar metallicity, POSYDON version 2 will include grids for 8 discrete metallicities: $2Z_{\odot}$, $Z_{\odot}$, $0.45Z_{\odot}$, $0.2Z_{\odot}$, $10^{-1}Z_{\odot}$, $10^{-2}Z_{\odot}$, $10^{-3}Z_{\odot}$, $10^{-4}Z_{\odot}$ (Andrews et al., 2024).
To evaluate the performance of interpolators trained on each of these MESA grids, we have three corresponding testing grids of binaries that are generated by random sampling of the same parameter spaces as the main binary grids. The end of MESA evolution in our models occurs when a star reaches the end of core carbon burning or the binary enters a common envelope phase (dynamically unstable mass transfer); more details about termination conditions can be found in Section 5.2 of Fragos et al. (2023). The outputs of these MESA grids include various properties of the binary at the end of MESA evolution, such as the binary orbital period, information about stellar structure, type of mass transfer experienced if any, and integrated or surface quantities of the stars, such as mass, radius, effective temperature, and luminosity. POSYDON also saves the stellar profiles at the end of MESA evolution, which show the structure of quantities like log density, log temperature, and hydrogen mass fraction at discrete sampling points throughout the star from the center to the surface as a function of either radius or mass. The placement of profile points is dictated by the adaptive mesh in the MESA model; points are more densely sampled in regions where higher resolution is necessary to characterize the complex behavior exhibited in the star (Paxton et al., 2011). The specific quantities saved after a MESA run can be specified by the user generating their own grids; the default settings are explained in detail in Fragos et al. (2023). As an example of the profiles generated by MESA, in Figure 1 we show the final log density profiles of the binaries in the CO-HMS testing grid.
@@ 2.1 Pre-processing
$\varepsilon_{i}=\left|\frac{\rho_{200}(m_{enc,i})-\rho_{400}(m_{enc,i})}{\rho_{400}(m_{enc,i})}\right|,$ (1)
@@ 3 Methods
While the fully-connected neural networks we use for emulation are less interpretable than, for example, tree-based methods, the high accuracy, scalability and computational efficiency of evaluation provide a good fit for our profile emulation task. Neural networks have previously been used by Maltsev et al. (2024) to efficiently emulate single-star evolutionary tracks calculated with MESA. This track emulation application, like the profile emulation presented here, is very useful for population synthesis studies. Neural network emulation has also been applied to many other astrophysical simulations, including by Ribas et al. (2020), who use neural networks to emulate synthetic protoplanetary spectral energy distributions at any point in their model’s parameter space. In this application, neural networks make it possible to efficiently sample posterior distributions for inference on protoplanetary disk masses, which would not be possible with the full radiative transport modeling. Auld et al. (2008) likewise use neural networks to emulate cosmic microwave background power spectra to ultimately conduct parameter estimation on cosmological parameters.
Gaussian process regression is sometimes used in emulation of astrophysical model outputs, for example observable stellar properties from single-star MESA simulations (Li et al., 2022), stellar spectra (Czekala et al., 2015), and cosmological simulations (Habib et al., 2007). However, Gaussian processes are not well suited for high-dimensional problems or large datasets due to the large amount of memory required in computing the covariance matrix (Rasmussen and Williams, 2006).
@@
To reduce our profiles’ dimensionality we use principal component analysis (PCA), which is a common technique in astrophysics. PCA provides a representation space in which data points are most widely distributed by choosing axes along which the dataset displays the most variance (Jolliffe, 2002). The first principal component is along the direction with maximal variance, the second principal component is orthogonal to the first along the direction that maximizes the residual variance in the dataset, and so on.
@@ 3.1 Emulating density profiles
In this context, the ‘profile’ refers to the array of values describing a given stellar property (e.g., density) as a function of the fraction of mass enclosed, and the ‘profile min’ and ‘profile max’ denote the minimum and maximum values of this array, respectively.
@@
Figure 2: Principal components of CO-HMS grid density profiles plus offsets for visual distinction. The first eight components are shown plotted in order of explained variance from top to bottom.
@@ 3.1.1 Dimensionality reduction
We first reduce the dimensionality of our normalized profiles to simplify the task assigned to the neural network. The smoothly-varying nature of stellar quantities like density and temperature make PCA a suitable method. In this process, each density profile, originally represented by 200 discrete data points sampled along the ‘fraction of mass enclosed’ axis (shown in Figure 1), is transformed into a lower-dimensional latent space. This latent space is spanned by a set of orthogonal 200-dimensional vectors, called principal components, which are derived from the training set and capture directions of maximum variance in the data. The first eight principal components of density profiles from the CO-HMS grid, which account for the most variance in those profiles, are shown in Figure 2. Each principal component contains 200 points along the same ‘fraction of mass enclosed’ axis as the original profiles. The y-axis scaling of the principal components is arbitrary, as they represent orthogonal directional vectors in the data space; for clarity, we have vertically offset them in the figure for visual distinction. Each training profile is projected onto the set of principal components to obtain the projection weights, which are used to create a reconstruction of the original profile as a weighted sum of the principal components.
@@ 3.1.2 Network Architecture
Our emulator is based on fully-connected feed-forward neural networks, with the input layer consisting of three input nodes (for each of the dimensions of the MESA grids, which are the two stellar masses and orbital period). The networks have 8 output nodes for the profiles’ principal component weights. The depth and width of the network were chosen to balance model complexity and prediction accuracy with computational efficiency. To do this, we conducted a grid search over various network sizes. The evaluation criteria for optimal performance were based on the percentage of median profile errors (explained in detail in Section 4) below 10% and below 100%. Among the configurations that performed best under these metrics, we further analyzed the full error distributions and selected the architecture that visually demonstrated the most desirable overall error distribution. The configuration of 12 hidden layers with 256 nodes per layer, shown in Figure 5, was determined to be the most effective architecture for our study, providing the best overall performance across the combined testing data from all mass transfer classes in all three binary grids. We experimented with including normalization layers (such as batch normalization) to stabilize learning and potentially improve convergence, but we found that this led to worse performance, potentially having interfered with the model’s ability to capture sufficient complexity. As a result, the normalization layers were excluded from the final architecture.
@@
We find that separating our data according to mass transfer class improves performance, so for each grid we produce one machine learning model for each mass transfer class. This is consistent with the existing initial-final interpolators from POSYDON v1 (Fragos et al., 2023). For training each of these profile emulation models, we used an 80/20 split to produce a training set and validation set. This step ensures the model’s ability to generalize outside the training data by identifying potential underfitting or overfitting. Each network was initialized randomly using the Glorot uniform kernel initializer and then given up to 3000 epochs to train to ensure the model had enough time to converge. However, to further prevent overtraining we used an “early stopping” feature to stop training if the validation loss does not improve for 100 epochs. This number, called the “patience”, was determined through experimentation to give the best results in optimization.
@@
The models were optimized based on a mean squared error (MSE) loss function placing equal emphasis on each component’s weight. It may seem more intuitive to base the loss function on the accuracy of profile predictions rather than the weights, because the accuracy of the weights is incidental to the ultimate goal of profile emulation. However, in practice we found that using a loss function based on the reconstructed profiles rather than the component weights ultimately made a minimal difference to the prediction accuracy. Hence, we chose the MSE loss of the weights for its faster evaluation time. We chose the Adam optimization algorithm, which is a stochastic gradient descent method that dynamically adjusts learning rates for each parameter based on the gradient (Kingma and Ba, 2014). This makes it suitable for a wide range of tasks, including those with sparse gradients or complex loss surfaces. We found that Adam converged to a significantly lower loss compared to other gradient descent algorithms such as SGD (Ruder, 2016), RMSprop (Hinton, 2012), AdamW (Loshchilov and Hutter, 2019), and Adagrad (Duchi et al., 2011). We also found that Adam provided faster convergence and more stable training for our dataset and model architecture.
@@
We use the ReLU (Rectified Linear Unit) activation function for all of the hidden layers, which allows the networks to learn non-linear behavior in the training set. In our experiments, ReLU outperformed other activation functions in model accuracy, including tanh (hyperbolic tangent) and sigmoid, for which the models had trouble converging during the optimization process. We found that tanh and sigmoid constrained layer outputs too tightly, which reduced the models’ ability to capture the complexity afforded by additional layers. By contrast, ReLU avoids these issues due to its simplicity and unbounded output for positive inputs, enabling the network to represent more complex patterns in the data. The output layer uses linear activation to ensure that final predictions for principal component weights are unconstrained. The model was implemented using TensorFlow (Abadi et al., 2015) using the Keras API (Chollet et al., 2015).
Figure 6: Flowchart showing procedure for emulating density profiles. The input binary conditions $M1,M2,P_{orb}$ represent the mass of the primary star, the mass of the secondary star, and the orbital period, respectively.
@@
Red boxes are model components from POSYDON’s existing scalar interpolation module, blue boxes are the model components presented in this work, and white boxes are data products. POSYDON’s scalar interpolation uses the initial parameters and associated MESA grids to predict final scalar properties at the end of an evolutionary stage. For density profiles, the final mass and central log density of each star is predicted by this interpolator; the emulator presented here then produces a density profile which is re-scaled by these predicted values.
@@
The overall procedure for obtaining predictions (illustrated in the flowchart in Figure 6) is as follows: we perform PCA on the training profiles to obtain a set of eight principal components and the eight principal component weights corresponding to each training profile. We then train neural networks with the training binaries’ initial conditions (star 1 mass, star 2 mass, orbital period) as input and eight principal component weights as outputs. With this network in place, we predict the weights for the profiles of a star in any new binary, which combined with the principal components construct predicted normalized profiles. We then renormalize these profiles using values for central log density and final star mass predicted with POSYDON’s scalar interpolation module, which predicts these quantities from the initial conditions and the associated MESA grid. With this configuration and the design choices detailed above, we achieve sufficient performance for profile emulation. Future work experimenting with different choices or architectures may further improve accuracy.